OpenStudy (anonymous):

Need help with rationalizing denominator (8√2)/(√20-√18)

OpenStudy (anonymous):

I got \[8\sqrt{40} +24\] but the answer is \[8\sqrt{10} +24\] and idk where i went wrong need help plz

OpenStudy (helder_edwin):

\[\large \frac{8\sqrt{2}}{\sqrt{20}-\sqrt{18}}\cdot \frac{\sqrt{20}+\sqrt{18}}{\sqrt{20}+\sqrt{18}}= \] \[\large =\frac{8\sqrt{2}(2\sqrt{5}+3\sqrt{2})}{20-18}= \frac{16\sqrt{10}+24(2)}{2}=8\sqrt{10}+24 \]

OpenStudy (anonymous):

how do you get \[16\sqrt{10} \]

OpenStudy (anonymous):

oh wait nvm lol